# Sub-shot noise sensitivity via deformed four-headed kitten states

### v2

@@ IV Enhancement in sensitivity
Sensitivity and its relationship to the phase-space characteristics of a quantum state are thoroughly discussed in the §II.1. This concept is then applied to the compass state in the §II.2, where it is demonstrated that the presence of sub-Planck structures in those states significantly have enhanced their sensitivity to displacement and that optimizing controlling parameters may further enhance this sensitivity far better than the standard quantum limit. In §III.4, we thoroughly examined the phase space of the proposed SA and AS cases, confirming the presence of sub-Planck structures in their phase spaces. We now examine how these sub-Planck structures have impact on the sensitivity to phase-space displacement, which is analyzed by assessing the sensitivities using Eq. (3).
First, for the SA scenario, we denote the associated sensitivity as $S_{\ket{\square}}(\delta)$, and is calculated as
@@
$O_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\delta):=$ $(\alpha_{i}\alpha^{*}_{j})^{q}G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{r}_{n=0}(-1)^{n}\Gamma(\alpha^{*}_{i}-\delta^{*})^{r-n}$
@@
Let us now examine the SA cases presented in the Figs. 11(a)-(c) belong to the SA situation, where it is evident that the overlap is zero for values $|\delta|<1$ (less than a coherent state) along arbitrary directions in phase space. This indicates that the sensitivity to displacement in this scenario surpasses the standard limit. The enhancement in sensitivity becomes more pronounced as the parameter $r$, which represents the number of added photons to the kitten state, increases. As shown in Figs. 11(a) and 11(b), this effect becomes clearly noticeable with a higher number of added photons. Specifically, the central structure is significantly reduced when the number of added photons increases from $r=12$ to $r=24$ with the number of subtracted photons, $q=12$, remains constant in this situation. Furthermore, the impact of increasing the number of eliminated photons from $q=12$ to $q=20$ is illustrated in the Figs. 11(b) and 11(c), where an enlargement in the central structure occurs, indicating that a higher value $|\delta|$ relative to the earlier case represented in Fig. 11(b) is needed to make the overlap zero. This suggests that the sensitivity to displacement in this circumstance decreases with an increment in $q$; hence, contrary to the parameter $r$, increasing the number of subtracted photons reduces sensitivity in the SA scenario.
In the AS scenario, as depicted in Figs. 11(d)-(f), comparable to the SA example, an improved sensitivity to phase space displacements $\delta$ is observed. The values $|\delta|<1$ can render the $S_{\ket{\blacksquare}}(\delta)$ zero as observable by the central structure. Similar to the SA example, this amplification becomes more noticeable as $r$ increases, as illustrated in the instances depicted in Figs. 11(d) and 11(e), where an increment in the $r$ as shown in Fig. 11(e) has reduced the central structure corresponding to the overlap function, meaning that now a smaller value of $|\delta|$ is required to make the overlap zero compared to the case shown in Fig. 11(d). Furthermore, we observe that in Fig. 11(e) and Fig. 11(f), the photon subtraction $q$ also does not affect the sensitivity enhancement as depreciated; in Fig. 11(f), raising $q$ results in the same overlap as its preceding instance. However, the photon subtraction $q$ may have an effect on the isotropic nature of the sensitivity, as optimal parameter selections may produce isotropic regions, such as the circle-type regions centered at the origin observed in Figs. 11(g) to 11(i) show that in the AS scenario, sensitivity is consistently increased in all directions, making these instances better compared to their counter parts of SA and compass states.
@@ V Outlook
We now provide a brief discussion on our results with their summarized physical consequences. This comprehensive discussion strives to integrate our findings into current understanding, providing a detailed picture of their effects and contributions to the field.
@@
A four-headed kitten state is considered a main example in our study, which represents a smaller version of a compass state [13], and our investigations and discussion are centered around this state, which serves as the foundation for our exploration and examination. The compass state exhibits fascinating sub-Planck scale structures and enhanced sensitivity. However, these characteristics are lost when transitioning to a kitten state [52], which is essentially a smaller version of the cat state, as the one we presented in Fig. 1(a). This transition occurs when the macroscopic parameter is reduced, resulting in a transformation from a cat state to a kitten state. Specifically, a bigger compass state exhibits pronounced sub-Planck features, indicating more finer quantum characteristics at scales lower than the Planck length, as discussed in §II.2 and then depicted in Fig. 1. Note that the mean photon number in bigger catlike states is higher [24], indicating enhanced total photon content and intensity. This comparison demonstrates the significant differences in quantum behavior and measurement precision between the larger compass and kitten states, as observed in Fig. 3.
@@
Photon addition and subtraction operations on squeezed-vacuum states are extremely useful approaches for creating larger cat states [50, 51, 52, 53]. These methods have also been demonstrated experimentally and provide an effective way to generate cat states of larger amplitude [52, 53]. In this study, we utilized the kitten version of the compass state and applied photon addition and subtraction operations with different order and magnitudes to construct our novel variants, as presented in Figs. 8, 9 and 10. We then investigated the phase-space characteristics of these variants to gain insights into their quantum properties. The kitten version of the compass state, as shown in Fig. 1(a), which does not possess sub-Planckness, is now transformed into the states holding sub-Planck structures and demonstrating an enhanced sensitivity, implying the effectiveness of these multi-photon processes, as evident in Fig. 11.
@@
When the number of photons added increases, nonclassical features contained by these states are improved, but increasing photon subtraction destroys nonclassical sub-Planck structures in phase space. Note that this only occurs for our SA case; otherwise, when photon subtraction is applied to a state directly (AS cases), the phase space nearly remains unchanged [see Fig. 5], and the size of the sub-Planck structures of the AS cases stays constant over the variation of the photon subtraction operations. Furthermore, photon addition raises the average photon count in the states, whereas photon subtraction maintains the photon statistics invariant. This demonstrates that the sequence of photon addition and subtraction operations in our case has a direct effect on the amount of quanta present in the states, which may be useful in maintaining the photon number in our states rather efficiently.
@@
Our quantum states exhibit comparable phase-space attributes as the compass state when there is a large number of the photon addition of the photons is applied to the four-headed kitten state, and the role of the photon subtraction in our instances is also interesting as in some cases it slightly enhances the isotropy of these features. Distinct phase-space characteristics between our SA and AS examples are highlighted, namely that AS cases have smaller sub-Planck structures and so achieve higher sensitivity when compared to their SA counterparts. Our investigation implies that variants of the four-headed kitten state we provided are an appropriate substitute for compass states and may perform better compared to the compass state under optimum conditions.
@@ VI Conclusion
We introduced alternative versions of the compass state, which are obtained by adding (and subtracting) an immense number of photons to a four-headed kitten state, with the option that the order in which these photons are applied to the state also changes and the noncommutativity of the bosonic operators results in two different quantum states with distinct phase-space features. Our outcomes revealed that the multiphoton operations we performed on the four-headed kitten state are quite effective and that these operations have transformed this multicomponent kitten state to other forms of states, which are richer in their nonclassical phase-space features and also exhibit sub-Planck structures. This investigation is relevant to previous experimental studies [52, 53], which demonstrated that photon operations are an effective means of enhancing the phase-space characteristics of quantum states, thereby supporting our findings.
The presence of crucial sub-Planck structures in the present investigation is influenced by the number of photons added or subtracted. Adding photons helps preserve these structures in both scenarios we explored, while subtracting photons typically disrupts them when photon addition is followed by photon subtraction. However, photon subtraction alone has no effect on the sub-Planck structures if applied before photon addition, and in this case (subtraction then addition), specific choices of photon subtraction operations may lead to isotropic versions of sub-Planck structures. These results directly apply to the sensitivity of these states as well. The sensitivity to displacement of the quantum states we proposed exceeds the standard limits, and this enhancement is also controlled by multiphoton operations. Specifically, increasing photon addition enhances sensitivity, while photon subtraction reduces it, in the sequence where photon addition is followed by subtraction. However, in the converse case of photon operations, the sensitivity to displacement remains unchanged over the variation of the photon subtraction, although, interestingly, for this case, improvement in the isotropy of sensitivity is observed for certain photon subtraction choices. The induction of sub-Planck structures in our indicated situations connects them to recent techniques for the development of nonclassical traits in quantum states [50, 51, 52, 53]. Perhaps these techniques may also be applied in the development of the quantum states we provided. Future research may inquire how to create our proposed quantum states, which will require a novel and thorough investigation to develop new techniques for their generation. This endeavor may involve formulating innovative strategies and methodologies specifically designed to produce these advanced quantum states.
@@ Acknowledgment
This work is supported by the Natural Science Foundation of Jiangsu Province (Grant No. BK20231320) and the National Natural Science Foundation of China (Grant No. 12174157).

### v3

@@ IV Enhancement in sensitivity
Sensitivity and its relationship to the phase-space characteristics of a quantum state were thoroughly discussed in the §II.1. These concepts were then applied to the compass state in §II.3, where it was demonstrated that the presence of sub-Planck structures in those states significantly enhances the sensitivity to displacement and that varying controlling parameters may further enhance this sensitivity far better than the standard quantum limit. In §III.4, we thoroughly examined the phase space of the proposed SA and AS cases of the four-headed kitten state, confirming the presence of sub-Planck structures in their phase spaces. We now examine how these sub-Planck structures have an impact on the sensitivity to phase-space displacement, which is analyzed by assessing the sensitivities using Eq. (3).
First, for the SA scenario, we denote the associated sensitivity by $S_{\ket{\square}}(\delta)$, which is calculated as
@@
$O_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\delta):=$ $(\alpha^{*}_{i}\alpha_{j})^{q}G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{r}_{n=0}(-1)^{n}\Gamma(\alpha^{*}_{i}-\delta^{*})^{r-n}$
@@
Let us now examine the SA cases presented in Figs. 11(a)-11(c), where it is evident that the overlap is zero for values $|\delta|<1$ (less than a coherent state) along arbitrary directions in phase space. This concept is also illustrated in detail in Appendix A, which reflects that the sensitivity to displacement in this scenario surpasses the standard limit. The enhancement in sensitivity becomes more pronounced as the parameter $r$, which represents the number of added photons to the four-headed kitten state, increases. As shown in Figs. 11(a) and 11(b), this effect becomes clearly noticeable with a higher number of photons added. Specifically, the central structure is significantly reduced when the number of added photons increases from $r=12$ to $r=24$ while the number of subtracted photons $q=12$ remains constant in this situation. Furthermore, the impact of increasing the number of eliminated photons from $q=12$ to $q=20$ is illustrated in the Figs. 11(b) and 11(c), where an enlargement in the central structure occurs, indicating that a higher value of $|\delta|$ relative to the earlier case represented in Fig. 11(b) is needed to make the overlap zero. This suggests that the sensitivity to displacement in this circumstance decreases with an increment in $q$; hence, contrary to the parameter $r$, increasing the number of subtracted photons reduces sensitivity in the SA scenario.
Figure 11: Overlap between a state and its slightly translated version for [(a)-(c)] SA examples and [(d)-(i)] comparable AS cases: (a) $r=12$ and $q=12$, (b) $r=24$ and $q=12$, (c) $r=24$ and $q=20$, (d) $r=12$ and $q=12$, (e) $r=24$ and $q=12$, (f) $r=24$ and $q=20$, (g) $r=q=1$, (h) $r=q=5$, and (i) $r=q=9$. In all situations, we set $c_{0}=1$ and normalize the intensity plots to unity.
In the AS scenario, as depicted in Figs. 11(d)-11(f), comparable to the SA example, an improved sensitivity to phase space displacements $\delta$ is observed. The values $|\delta|<1$ can render the $S_{\ket{\blacksquare}}(\delta)$ zero as observable by the central structure. Similar to the SA example, this amplification becomes more noticeable as $r$ increases, as illustrated in the instances depicted in Figs. 11(d) and 11(e), where an increment in $r$ as shown in Fig. 11(e) has reduced the central structure corresponding to the overlap function, meaning that now a smaller value of $|\delta|$ is required to make the overlap zero compared to the case shown in Fig. 11(d). Furthermore, we observe that in Figs. 11(e) and 11(f), the photon subtraction $q$ does not reduce the sensitivity enhancement, increasing $q$ results in the same overlap as its preceding instance. However, the photon subtraction $q$ may have an effect on the isotropic nature of the sensitivity, as optimal parameter selections may produce isotropic regions. For example, the circle-type regions centered at the origin observed in Figs. 11(g)-11(i) show that in the AS scenario, sensitivity is consistently increased in all directions, making these instances better compared to their counter-parts of SA and compass states.
@@ V Outlook
We now provide a brief discussion of our results with their summarized physical consequences. This comprehensive discussion strives to integrate our findings into current understanding, providing a detailed picture of their effects and contributions to the field.
@@
A four-headed kitten state is considered as main example in our work, which represents a smaller version of a compass state [13], and our investigation and discussion are centered around this state, which serves as the foundation for our exploration and examination. The compass state exhibits fascinating sub-Planck scale structures and enhanced sensitivity. However, these characteristics are lost when transitioning to a kitten state [75], which is essentially a smaller version of the cat state, such as the one we presented in Fig. 1(a). This transition occurs when the macroscopic parameter is reduced, resulting in a transformation from a cat state to a kitten state. Specifically, a bigger compass state exhibits pronounced sub-Planck features, indicating more finer quantum characteristics at scales lower than the Planck length, as discussed in §II.3 and then depicted in Fig. 1. Note that the mean photon number in bigger catlike states is higher [36], indicating enhanced total photon content and intensity. This comparison demonstrates the significant differences in quantum behavior and measurement precision between the larger compass and kitten states, as observed in Fig. 3.
@@
Photon addition and subtraction operations on squeezed-vacuum states are extremely useful approaches for creating larger cat states [73, 74, 75, 76]. These methods have also been demonstrated experimentally and provide an effective way to generate cat states of larger amplitude [75, 76]. In this work, we utilized the kitten version of the compass state and applied photon addition and subtraction operations with different order and magnitudes to construct our variants, as presented in Figs. 8, 9 and 10. We then investigated the phase-space characteristics of these variants to gain insights into their quantum properties. The kitten version of the compass state, as shown in Fig. 1(a), which does not possess sub-Planckian scale, is now transformed into the states holding sub-Planck structures and demonstrating an enhanced sensitivity, implying the effectiveness of these multiphoton processes, as evident in Fig. 11.
@@
When the number of photons added increases, nonclassical features contained by these states are improved, but increasing photon subtraction destroys nonclassical sub-Planck structures in phase space. Note that this only occurs for our SA case; otherwise, when photon subtraction is applied to a state directly (AS cases), the phase space nearly remains unchanged [see Fig. 5], and the size of the sub-Planck structures of the AS cases stays constant over the variation of the photon subtraction operations, but in this case, under the particular selection of the photon subtraction the isotropy of corresponding sub-Planck structures is improved, as observed in Fig. 10(a)-10(c). Furthermore, photon addition raises the average photon count in the coherent states, whereas photon subtraction keeps the photon statistics invariant. This demonstrates that the sequence of photon addition and subtraction operations in our case may have a direct effect on the amount of quantum characteristics present in the states.
@@
Our quantum states exhibit phase-space attributes comparable to the compass state when there is a large number of photon applied to the four-headed kitten state. In addition, the role of photon subtraction in our instances is also interesting as in some cases it greatly enhances the isotropy of these features. Distinct phase-space characteristics between our SA and AS examples are highlighted, namely, that AS cases have smaller sub-Planck structures and so achieve higher sensitivity when compared to their SA counterparts. Our investigation implies that variants of the four-headed kitten state we provided are an appropriate substitute for compass states and may perform better compared to the compass state under optimum conditions.
@@ VI Conclusion
We have introduced alternative versions of the compass state, which are obtained by adding (and subtracting) an immense number of photons to a four-headed kitten state, with the option that the order in which these photons are applied to the state also changes and the noncommutativity of the bosonic operators results in two different quantum states with distinct phase-space features. Our outcomes revealed that the multiphoton operations we performed on the four-headed kitten state are quite effective and that these operations have transformed this multicomponent kitten state to other forms of states, which are richer in their nonclassical phase-space features and also exhibit sub-Planck structures. Our investigation also has a close connection with previous experimental studies [75, 76], which demonstrated that photon operations are an effective means of enhancing the phase-space characteristics of quantum states, thereby supporting our findings.
The presence of crucial sub-Planck structures in the present investigation is influenced by the number of photons added or subtracted. Adding photons helps preserve these structures in both scenarios we explored, while subtracting photons typically disrupts them when photon addition is followed by photon subtraction. However, photon subtraction alone has no effect on the sub-Planck structures if applied before photon addition, and in this case (subtraction then addition), specific choices of photon subtraction operations may lead to isotropic versions of sub-Planck structures. These results directly apply to the sensitivity of these states as well. The sensitivity to displacement of the quantum states we proposed exceeds the standard limits, and this enhancement is also controlled by multiphoton operations. Specifically, increasing photon addition enhances sensitivity, while photon subtraction reduces it, in the sequence where photon addition is followed by subtraction. However, in the converse case of photon operations, the sensitivity to displacement remains unchanged over the variation of the photon subtraction, although, interestingly, for this case, improvement in the isotropy of sensitivity is observed for certain photon subtraction choices. The induction of sub-Planck structures in our indicated situations connects them to recent techniques for the development of nonclassical traits in quantum states [73, 74, 75, 76], and perhaps these techniques further may also be applied in the development of the quantum states we provided. Future research may inquire how to create our proposed quantum states, which will require a novel and thorough investigation to develop new techniques for their generation. This endeavor may involve formulating innovative strategies and methodologies specifically designed to produce these advanced quantum states.
@@ Acknowledgment
This work was supported by the Natural Science Foundation of Jiangsu Province (Grant No. BK20231320) and the National Natural Science Foundation of China (Grant No. 12174157). We would like to thank Tan Hailin for the valuable suggestions and comments on the work.
Appendix A sub-Planckian scale
In this appendix, we provide an overview of the fundamental concept of the sub-Planck structure, utilizing graphical illustrations to effectively explore and elucidate this concept. These visual illustrations help in providing a clearer understanding of the intricate details and underlying principles associated with sub-Planck scales.
Figure 12: [(a)-(h)] Wigner distributions for (a) vacuum state ($\alpha=0$), (b) the kitten state with $c_{0}=1$, (c) the central interference of the horizontal cat state with $c_{0}=5$, (d) the four-headed kitten state of Fig. 1(a), and the cases corresponding to (e) Fig. 1(b), (f) Fig. 9(a), (g) Fig. 9(d), and (h) Fig. 10(c), which are now plotted with the regions where these distributions get minor amplitudes highlighted by black lines, which are enclosing the corresponding phase-space features. Also shown are the overlap functions for different quantum states of the main text: (i) Overlap between a horizontal cat and its displaced version and [(j)-(l)] overlap functions corresponding to the cases presented in (j) Fig. 3(a), (k) Fig. 11(a), and (l) Fig. 11(i) with white lines marking the zeros of each function.
Planck scale. The Wigner function of the coherent state, as given by Eq. (4), has the phase-space dimension at the Planck scale; that is, the coherent state follows the equality of the Heisenberg uncertainty principle and is considered the minimum uncertainity state [16]. The Wigner function shown in Fig. 12(a) for $\alpha=0$ (a vacuum state) represents a phase-space distribution where the uncertainty in position $\Delta x$ and momentum $\Delta p$ is as small as that set by the minimal standard quantum limit. This means the ‘volume’ of phase space, given by the product $\Delta x\Delta p$, is at its smallest possible value ($\Delta x\Delta p\sim\nicefrac{{1}}{{2}}$), in line with the Heisenberg uncertainty principle. Additionally, the spread of the Wigner function in phase space stays the same, meaning its overall shape remains fixed over the variation of associated parameters, staying within the boundaries set by the uncertainty constraint. Hence, simply the phase-space volume of the coherent sets the Planck scale, and the phase-space features of extensions below this norm are simply called sub-Planck scale features [13].
The smallest phase-space feature associated with the four-headed kitten state investigated in this work also adheres to the Heisenberg uncertainty principle. For example, the case described in the Fig. 1(a) is replotted with the comparison of the coherent state in Fig. 12(d) and appears to demonstrate a phase-space volume comparable to that of the coherent state and is consequently considered to be at the Planck scale level.
Sub-Planckian scale. Figures  12(c) and 12(e) depict the central phase-space areas of the cat and compass states, respectively, as explained in the main text. The illustrations emphasize the zeros of the relevant distributions, demonstrating the limitations of corresponding features in the anisotropic or isotropic domains. The black solid lines indicate areas where the Wigner function achieves amplitudes of $10^{-2}$. Upon closer inspection of the central features within this distribution, we notice that they are confined along the black solid lines, which trace out a tile-like pattern in phase space. It is immediately apparent that the size of this central phase-space region is significantly smaller than that of the coherent state depicted in Fig. 12(a), exhibiting a form of anisotropic sub-Planck structures. Similar sub-Planck structures are present in the situations appearing in Figs. 12(f), 12(g), and 12(h), which belong to the situations shown in Figs. 9(a), 9(d), and 10(c), respectively. In Fig. 10(l) the core sub-Planck structures are contained in a circular region, indicating an isotropic version of the structure.
Sensitivities. The overlap between the horizontal cat state and its slightly displaced constituent is depicted in Fig. 12(i). The white lines enclosing structures underline the corresponding zeros of the distributions. The overlap for the compass state case, as shown in Fig. 3(a), is now represented with the zeros allocated along white lines in Fig. 12(j). This makes it evident that the central structure has a volume smaller than that of the coherent state, for which $|\delta|<1$ is needed to make the overlap zero. In addition, anisotropic enhancement in the sensitivity is observed, leading to the similar cases shown in Figs. 12(k) and 12(l), which correspond to the cases illustrated in the main text in Figs. 11(a) and 11(i) respectively. The case presented in Fig. 11(i) represents the situation in which isotropic enhancement in the corresponding sensitivity is observed.